# Explainable Machine Learning-Based Security and Privacy Protection Framework for Internet of Medical Things Systems

### v1

@@ 4.2 Dataset Description
Experiments were carried out on four distinct datasets containing attacks that could potentially impact the availability, confidentiality, and integrity of IoMT systems. These datasets include NSL-KDD, UNSW-NB15, and ToN-IoT, all comprising network data, and WUSTL-EHMS, which encompasses both network and medical data. Each dataset is described in detail as follows:
@@
  • WUSTL-EHMS : Using a real-time Enhanced Healthcare Monitoring System (EHMS) testbed, the WUSTL-EHMS dataset was generated [15]. Due to the limited availability of a dataset that integrates these biometrics, this testbed accrues both the network flow metrics and patients’ biometrics. This dataset comprises man-in-the-middle attacks, spoofing and data injection. The spoofing attack merely sniffs the protocols passing through the gateway and the server, thus infringing upon the confidentiality of the patient’s data. The data injection attack is employed to dynamically modify the packets, thereby infringing upon the integrity of the data.
@@ 4.3 Data preprocessing
During the data preprocessing phase, a diverse array of methods was employed to optimize the dataset for analysis. To begin, a mapping function was utilized to assign binary values, effectively distinguishing instances representing attacks from those that did not, thereby enabling a clear classification of the data. The transformation of Boolean values into binary further streamlined the dataset’s representation. For categorical features, the ordinal and OneHotEncoder techniques were applied, ensuring their proper handling in the subsequent analysis. In order to maintain an unbiased approach, certain features primarily used for dataset labeling were meticulously removed. Additionally, features that held a single value, bearing no meaningful information for the model’s learning process, were excluded from the dataset. Subsequently, numerical features underwent standardization using the StandardScaler method, optimizing their scale and improving their compatibility with the model.
Ultimately, the dataset was partitioned into an 80% training set and a 20% testing set. In the context of FL, the 80% of the training data are independently and identically distributed among clients, while the remaining 20% of the test set is utilized for evaluating the global model. This division facilitated the model’s learning process on the training set, while the testing set provided an independent evaluation of the model’s performance. These preprocessing steps were essential in preparing the data for effective model training and evaluation.
@@ 4.4 ML Algorithm
The construction of various ANN models tailored to specific datasets involved distinctive configurations. For the UNSW-NB15 dataset, ANNs were constructed with seven hidden layers, each comprising a different number of units: 150, 120, 90, 60, 30, 20, and 10 units, respectively. Conversely, the ToN-IoT dataset utilized an ANN model with five hidden layers, incorporating 60, 40, 30, 20, and 10 units, respectively. In the case of the NSL-KDD dataset, a five-layered ANN model was established, featuring 80, 40, 30, 20, and 10 units. An ANN model was constructed for the WUSTL-EHMS dataset, comprising three hidden layers that were specified to contain 10, 20, and 40 units, respectively.
@@ 4.5 Evaluation metrics
In the following formulas, True Positive (TP) denotes instances that were correctly identified as positive, False Negative (FN) represents instances that were incorrectly classified as negative, False Positive (FP) stands for instances incorrectly classified as positive, and True Negative (TN) corresponds to instances correctly classified as negative. These measures are indispensable for assessing the overall performance of a classification model [55].
  • Confusion matrix : The confusion matrix is a tabular representation that provides a summary of the performance of a classification model through the counts of TP, TN, FP and FN [55].
@@ 4.6 Experiment Results
The testing methodology encompassed a meticulous examination of several crucial elements. First and foremost, an emphasis was placed on the selection of optimal parameters within the FL process. This involved a systematic exploration to attain peak performance, with a focus on achieving high rates for each tested metric while concurrently reducing the number of communication rounds. Key aspects considered in this optimization process included increasing the number of participating clients, elevating the fraction of clients engaged in each FL round, and boosting the local epochs executed by participating clients.
Furthermore, the study involved a meticulous comparison between the results derived from the optimized FL approach, showcasing superior performance, and those obtained from a centralized IDS. This comparative analysis aimed to provide valuable insights into the efficacy of the FL model in relation to the traditional centralized approach.
To enhance the interpretability and understanding of the different models under examination, the SHAPE method was applied. This post-analysis method allowed for the interpretation and explanation of the top-performing models across various datasets. The incorporation of SHAPE into the evaluation process aimed to offer a deeper understanding of the intricacies and decision-making processes within the models that demonstrated exceptional performance across diverse testing scenarios.
metric / clients number 2 4 8
accuracy 0.9807 0.9808 0.9804
precision 0.9647 0.9618 0.9633
recall 0.9807 0.9841 0.9814
f1-score 0.9726 0.9728 0.9723
TP 31614 31724 31635
TN 58816 58712 58769
FP 1157 1261 1204
FN 622 512 601
loss 0.0521 0.0543 0.0524
auc 0.9978 0.9977 0.9976
communication rounds 52 56 48
@@
metric / clients number 2 4 8
accuracy 0.9881 0.9883 0.9881
precision 0.9431 0.9424 0.9511
recall 0.9482 0.9508 0.9392
f1-score 0.9456 0.9466 0.9451
TP 1061 1064 1051
TN 9083 9082 9093
FP 64 65 54
FN 58 55 68
loss 0.1009 0.0346 0.0542
auc 0.9982 0.9984 0.9983
communication rounds 80 33 37
@@
metric / clients number 2 4 8
accuracy 0.9856 0.9904 0.9905
precision 0.9883 0.99 0.9904
recall 0.9817 0.99 0.9898
f1-score 0.985 0.99 0.99
TP 13990 14108 14106
TN 15286 15309 15315
FP 166 143 137
FN 261 143 145
loss 0.237 0.0822 0.0652
auc 0.9846 0.9958 0.9973
communication rounds 50 27 18
@@
metric / clients number 2 4 8
accuracy 0.9378 0.9381 0.9381
precision 0.9339 0.9381 0.9626
recall 0.53 0.53 0.515
f1-score 0.6762 0.6773 0.671
TP 212 212 206
TN 2849 2850 2856
FP 15 14 8
FN 188 188 194
loss 0.1973 0.2 0.1904
auc 0.8927 0.8996 0.8997
communication rounds 15 16 24
@@
accuracy 0.9882 0.9882 0.9881
precision 0.9471 0.944 0.9511
recall 0.9446 0.9482 0.9392
f1-score 0.9459 0.9461 0.9451
TP 1057 1061 1051
TN 9088 9084 9093
FP 59 63 54
FN 62 58 68
loss 0.1697 0.208 0.0542
auc 0.9982 0.9982 0.9983
communication rounds 73 83 37
@@
accuracy 0.9887 0.9888 0.989 0.9886
precision 0.9448 0.9354 0.9498 0.953
recall 0.95 0.9618 0.9472 0.9399
f1-score 9474 0.9484 0.9485 0.9464
TP 1044 1057 1041 1033
TN 9106 9094 9112 9116
FP 61 73 55 51
FN 55 42 58 66
loss 0.0393 0.0261 0.0251 0.0281
auc 0.9989 0.999 0.999 0.999
communication rounds 30 12 5 2
@@ 4.6.1 FL Parameters Selections
The exploration of the dynamics of FL involved a meticulous analysis of various parameters that could impact performance and the number of communication cycles. These parameters include the number of patients, the fraction fit, and the number of local epochs.
To initially assess the impact of increasing the number of patients and participants in the FL process on performance and the number of communication cycles, the local epochs were set to 1, the fraction fit to 1, and the number of clients was varied to 2, 4, and 8.
The increase in the number of clients led to a noticeable reduction in communication cycles to achieve the desired performance levels. This trend was observed in the ton_iot and NSL-KDD datasets, achieving the lowest number of communication cycles with 8 clients compared to other tests, as shown in Figures 1(a) and 1(c) respectively. However, with the UNSW_NB15 dataset, the lowest number of communication cycles to achieve the desired performance was reached with 4 clients, close to that obtained with 8 clients but reducing more than half the number of communication cycles compared to using only 2 clients, as observed in Figure 1(b), confirming the trend observed with the other datasets, ton_iot and NSL-KDD. However, in the context of the WUSTL-EHMS dataset, an opposite trend is observed where increasing the number of clients implies an increase in the number of communication rounds to achieve the desired performance, as shown in Figure 1(d). This can be explained by the low number of samples in the dataset, leading to a limited availability of samples between clients, requiring more communication rounds to generalize.
Next, the examination shifts to the test of fraction fit and its impact on performance and the number of communication cycles. For this purpose, local epochs are set to 1, and the number of clients is fixed at 8, as demonstrated previously to reduce the number of communication rounds. Subsequently, the fraction fit percentage is varied to 0.1, 0.5, and 1.
Moreover, increasing the adjustment of the fit fractions proved effective in reducing the number of communication cycles while achieving the desired results. This is observed in the ton_iot, UNSW_NB15, and NSL-KDD datasets, illustrated in Figures 2(a), 2(b), and 2(c) respectively. However, for the WUSTL-EHMS dataset, the lowest number of communication rounds to achieve the desired performance is reached with a fraction fit of 0.5, slightly increasing in the case of fraction fit = 1, confirming the trend observed with the other datasets.
Finally, the examination focuses on the test of the number of local epochs and its influence on performance and the number of communication cycles. For this analysis, fraction fit is set to 1, and the number of clients is fixed at 8, given that these parameters have been demonstrated previously to reduce the number of communication rounds. Subsequently, the number of local epochs is varied to 1, 2, 5, and 8.
A remarkable observation was made regarding the increase in the number of local epochs, revealing a consistent reduction in communication cycles while achieving commendable performance. This observed trend was consistent for all tested datasets, as shown in Figures 3(a), 3(b), and 3(d), except for the NSL-KDD dataset where the best number of communication rounds is achieved with only 2 epochs, as shown in Figure 3(c), although it remains close to that obtained with 8 epochs, confirming the downward trend of increasing the number of local epochs on the number of communication rounds.
@@ 4.6.2 XAI results
Previous test results demonstrate the IDS system’s commendable performance in anomaly detection. However, an investigation into why the proposed solution predicts as it does is imperative. To address this concern, the SHAP method was employed, enabling the identification of feature relevance in anomaly detection. The outcomes of SHAP method can be depicted graphically, utilizing variable-length bars and color coding called beeswarm plot. This visualization effectively demonstrates how each level or range of values of a particular feature positively or negatively influences the classification result. Each point on the graph represents a feature value. Red points denote higher feature values, while blue points indicate lower feature values. Values on the left side of the x-axis tend toward the normal class, whereas those on the right side tend toward the anomaly class [60]. There is also other types of graphs, such as bar plots. In these plots, the x-axis represents the Shapley value, and the y-axis represents the feature names. Features with the most significant impact are positioned at the top of the graph, while those with the least impact are at the bottom [61].
@@
In this approach, the SHAP value is applied to the final FL model, achieving the best performance using the test dataset. This process produced two types of graphics: beeswarm plots and bar plots.
@@
The SHAP results for the UNSW-NB15 dataset are depicted in Figure 4, while the key features exerting the most significant impact on the anomaly detection classification process, as shown in Figure 3(b), include sttl (Time to live from the source to the destination), ct_state_ttl (Connection state value of time to live), and dttl (Time to live from the destination to the source). Notably, high values of these features, all linked to Time To Live (TTL), play a pivotal role in anomaly detection, visually represented in the accompanying figures 3(a). Prominent patterns emerge, underscoring that elevated values in sttl, ct_state_ttl, and dttl serve as indicators of anomalies, especially concerning TTL. Under normal circumstances, TTL values tend to exhibit stability or fall within specific ranges. However, higher values may signify abnormal extensions of connections, potentially employed to circumvent temporary security mechanisms. It is essential to note that certain attacks comprise very few samples, posing a challenge for ML in discerning crucial features for the detection of such anomalies.
@@
In the ToN-IoT dataset, as illustrated in Figure 5, the features wielding the most significant impact on anomaly detection classification include proto _udp (indicating the use of UDP), dst_port (representing destination ports), conn_state_oth (denoting other unspecified connection states), service_dns (reflecting the use of DNS services), conn_state_SF (indicating an established connection with successful data exchange), conn_state_rej (signifying connection rejection or inability to establish), and src_port (depicting source ports from the endpoint’s TCP/UDP ports), as shown in Figure 4(b). The absence of UDP protocol usage, as demonstrated in Figure 4(a), is pivotal in detecting anomalies, emphasizing its role in discerning abnormal network behavior. Notably, lower destination port numbers indicate ongoing attacks, with attackers often targeting well-known services associated with such ports (e.g., FTP=21, SSH=22, or HTTP=80). Correlated attributes, namely conn_state_oth, conn_state_SF, and conn_state_rej, play a critical role in attack detection. The absence of a specific connection state in conn_state_oth, as depicted in Figure 4(a), suggests potential interception or falsification of traffic, while conn_state_SF points to established connections without data exchange, possibly indicating suspicious activities. Simultaneously, conn_state_rej highlights repeated or massive attempts to establish rejected connections, hinting at DDoS attacks or vulnerability exploitation. The absence of DNS port usage, as shown in Figure 4(a), indicates a potential attack, as attackers may manipulate DNS requests to mask their activities. Additionally, elevated values of the source port attribute enhance attack detection, as attackers often initiate attacks from dynamic ports not assigned to well-known services like FTP.
@@
The SHAP results for the WUSTL-EHMS dataset, as illustrated in Figure 7, highlight features influencing the classification process in anomaly detection. These features, outlined in Figure 6(b), include Dintpkt (Destination Inter Packet), dstjitter (Destination Jitter), dstload (Destination Load), load, and srcload (Source Load). Each of these features holds distinct relevance for understanding and characterizing network behavior, especially in the context of health-related host data. Significantly, features associated with health data have a profound impact on detecting attacks, particularly those related to data injection. Anomalies in health-related metrics can be identified by values deviating from typical ranges. For example, elevated pulse_rate or unusually low temperatures, as depicted in Figures 6(a), can signal abnormal conditions. Elevated values of Dintpkt or dstjitter, which are correlated features, may indicate unusual intervals or abnormal temporal variations between packets or data received by the destination. In the context of spoofing, such variations could signal an attempt to manipulate network traffic, concealing or altering the true origin of packets. This observation aligns with the understanding that packets in spoofing scenarios are often manually modified by attackers, sent individually with extended time intervals between them, rather than in an organized flow. Low values of dstload, load, and srcload also play a significant role in anomaly detection. A destination with low load may be unusual in a normal context, potentially indicating an attacker not targeting system availability but engaging in other types of attacks, such as injection or scanning attacks. This behavior might be explained by the attacker’s intent to remain inconspicuous and avoid detection, highlighting the multifaceted role these features play in discerning anomalies within network behavior.

### v2

@@ 4.2 Dataset Description
Experiments are carried out on four distinct datasets containing attacks that can potentially impact the availability, confidentiality, and integrity of IoMT systems. These datasets include NSL-KDD, UNSW-NB15, and ToN-IoT, all comprising network data, and WUSTL-EHMS, which encompasses both network and medical data. Each dataset is described in detail as follows:
@@
  • WUSTL-EHMS : Using a real-time Enhanced Healthcare Monitoring System (EHMS) testbed, the WUSTL-EHMS dataset was generated [17]. Due to the limited availability of a dataset that integrates these biometrics, this testbed accrues both the network flow metrics and patients’ biometrics. This dataset comprises MITM attacks, spoofing and data injection. The spoofing attack merely sniffs the protocols passing through the gateway and the server, thus infringing upon the confidentiality of the patient’s data. The data injection attack is employed to dynamically modify the packets, thereby infringing upon the integrity of the data.
@@ 4.3 Data Preprocessing
During the data preprocessing phase, a diverse array of methods is employed to optimize the dataset for analysis. To begin, a mapping function is utilized to assign binary values, effectively distinguishing instances representing attacks from those that do not, thereby enabling a clear classification of the data. The transformation of Boolean values into binary further streamlines the dataset’s representation. For categorical features, the ordinal and OneHotEncoder techniques are applied, ensuring their proper handling in the subsequent analysis. To maintain an unbiased approach, certain features primarily used for dataset labeling are meticulously removed. Additionally, features that hold a single value, bearing no meaningful information for the model’s learning process, are excluded from the dataset. Subsequently, numerical features undergo standardization using the StandardScaler method, optimizing their scale and improving their compatibility with the model.
Ultimately, the dataset is partitioned into an 80% training set and a 20% testing set. In the context of FL, the 80% of the training data are independently and identically distributed among clients, while the remaining 20% of the test set is utilized for evaluating the global model. This division facilitates the model’s learning process on the training set, while the testing set provides an independent evaluation of the model’s performance. These preprocessing steps are essential in preparing the data for effective model training and evaluation.
@@ 4.4 ML Algorithm
The construction of various ANN models tailored to specific datasets involves distinctive configurations. For the UNSW-NB15 dataset, ANNs are constructed with seven hidden layers, each comprising a different number of units: 150, 120, 90, 60, 30, 20, and 10 units, respectively. Conversely, the ToN-IoT dataset utilizes an ANN model with five hidden layers, incorporating 60, 40, 30, 20, and 10 units, respectively. In the case of the NSL-KDD dataset, a five-layered ANN model is established, featuring 80, 40, 30, 20, and 10 units. An ANN model is constructed for the WUSTL-EHMS dataset, comprising three hidden layers that are specified to contain 10, 20, and 40 units, respectively.
@@ 4.5 Evaluation Metrics
In the following formulas, True Positive (TP) denotes instances that are correctly identified as positive, False Negative (FN) represents instances that are incorrectly classified as negative, False Positive (FP) corresponds to instances that are incorrectly classified as positive, and True Negative (TN) refers to instances that are correctly classified as negative. These measures are indispensable for assessing the overall performance of a classification model [57].
  • Confusion matrix : The confusion matrix is a tabular representation that provides a summary of the performance of a classification model through the counts of TP, TN, FP, and FN [57].
@@
  • AUC (Area Under the Curve) : is a metric used to evaluate the performance of a binary classification model, specifically by measuring the area under the ROC (Receiver Operating Characteristic) curve. It quantifies the model’s ability to distinguish between positive and negative classes, with values ranging from 0 to 1. An AUC of 1 indicates perfect classification, while 0.5 suggests no discriminative power, equivalent to random guessing [62].
@@ 4.6 Experiment Results
The testing methodology encompasses a meticulous examination of several crucial elements. First and foremost, emphasis is placed on the selection of optimal parameters within the FL process. This involves a systematic exploration to attain peak performance, with a focus on achieving high rates for each tested metric while concurrently reducing the number of communication rounds. Key aspects considered in this optimization process include increasing the number of participating clients, elevating the fraction of clients engaged in each FL round, and boosting the local epochs executed by participating clients.
metric / clients number 2 4 8 12
accuracy 0.9807 0.9808 0.9804 0.9792
precision 0.9647 0.9618 0.9633 0.9624
recall 0.9807 0.9841 0.9814 0.9787
f1-score 0.9726 0.9728 0.9723 0.9705
TP 31614 31724 31635 31614
TN 58816 58712 58769 58637
FP 1157 1261 1204 1303
FN 622 512 601 622
loss 0.0521 0.0543 0.0524 0.0576
auc 0.9978 0.9977 0.9976 0.9967
communication rounds 52 56 48 47
@@
metric / clients number 2 4 8 12
accuracy 0.9881 0.9883 0.9881 0.9884
precision 0.9431 0.9424 0.9511 0.9511
recall 0.9482 0.9508 0.9392 0.9398
f1-score 0.9456 0.9466 0.9451 0.9454
TP 1061 1064 1051 1031
TN 9083 9082 9093 9116
FP 64 65 54 53
FN 58 55 68 66
loss 0.1009 0.0346 0.0542 0.0266
auc 0.9982 0.9984 0.9983 0.9989
communication rounds 80 33 37 9
@@
metric / clients number 2 4 8 12
accuracy 0.9856 0.9904 0.9905 0.9904
precision 0.9883 0.99 0.9904 0.9927
recall 0.9817 0.99 0.9898 0.9873
f1-score 0.985 0.99 0.99 0.99
TP 13990 14108 14106 14036
TN 15286 15309 15315 15383
FP 166 143 137 103
FN 261 143 145 181
loss 0.237 0.0822 0.0652 0.0632
auc 0.9846 0.9958 0.9973 0.9971
communication rounds 50 27 18 17
@@
metric / clients number 2 4 8 12
accuracy 0.9378 0.9381 0.9381 0.9381
precision 0.9339 0.9381 0.9626 0.9157
recall 0.53 0.53 0.515 0.5704
f1-score 0.6762 0.6773 0.671 0.7029
TP 212 212 206 239
TN 2849 2850 2856 2823
FP 15 14 8 22
FN 188 188 194 180
loss 0.1973 0.2 0.1904 0.2061
auc 0.8927 0.8996 0.8997 0.9166
communication rounds 15 16 24 37
@@
accuracy 0.9892 0.9884 0.9884
precision 0.9482 0.9587 0.9511
recall 0.9508 0.9316 0.9398
f1-score 0.9495 0.945 0.9454
TP 1043 1022 1031
TN 9112 9125 9116
FP 57 44 53
FN 54 75 66
loss 0.0274 0.208 0.0266
auc 0.9989 0.999 0.9989
communication rounds 12 8 9
@@
accuracy 0.9882 0.9881 0.989 0.9888
precision 0.9516 0.943 0.9408 0.948
recall 0.9425 0.9512 0.9556 0.9521
f1-score 0.9470 0.9471 0.9481 0.95
TP 1082 1092 1097 1093
TN 9063 9052 9049 9058
FP 55 66 69 60
FN 66 56 51 55
loss 0.1047 0.0476 0.0301 0.0341
auc 0.998 0.998 0.9985 0.9984
communication rounds 25 10 5 3
@@
Furthermore, the study involves a meticulous comparison between the results derived from the optimized FL approach, showcasing superior performance, and those obtained from a centralized IDS. This comparative analysis aims to provide valuable insights into the efficacy of the FL model in relation to the traditional centralized approach.
To enhance the interpretability and understanding of the different models under examination, the SHAP method is applied. This post-analysis method allows for the interpretation and explanation of the top-performing models across various datasets. The incorporation of SHAP into the evaluation process aims to offer a deeper understanding of the intricacies and decision-making processes within the models that demonstrate exceptional performance across diverse testing scenarios.
@@ 4.6.1 FL Parameters Selections
The exploration of FL dynamics involves a systematic analysis of key parameters influencing performance and communication cycles. These parameters include the number of participating clients, the fraction fit, and the number of local training epochs.
To evaluate the impact of varying the number of clients on performance and communication cycles, the local epochs are fixed at 1, the fraction fit at 1, and the number of clients is adjusted to 2, 4, 8, and 12. Increasing the number of clients results in a significant reduction in communication cycles required to achieve target performance levels. This trend is evident in the Ton-IoT and NSL-KDD datasets, where the lowest number of communication cycles is achieved with 12 clients, and no significant variation is observed with 8 clients compared to other configurations, as shown in Tables 3(a) and 3(c). This suggests that communication cycles stabilize with 8 or more clients. For the UNSW_NB15 dataset, the lowest number of communication cycles is achieved with 12 clients, showing a significant reduction compared to 8 clients, as indicated in Table 3(b). This aligns with the decreasing trend in communication cycles as the number of clients increases, consistent with observations from the Ton-IoT and NSL-KDD datasets. In contrast, the WUSTL-EHMS dataset exhibits an inverse trend, where increasing the number of clients leads to a higher number of communication rounds to achieve target performance, as depicted in Table 3(d). This behavior is likely due to the limited sample size in the dataset, resulting in insufficient data distribution across clients and requiring additional communication rounds for model generalization.
Next, the analysis examines the influence of the fraction fit parameter on performance and communication cycles. For this evaluation, the number of local epochs is set to 1, and the number of clients is fixed at 8 for the Ton-IoT and NSL-KDD datasets, as communication rounds decrease and stabilize at this client count. Similarly, the WUSTL-EHMS dataset is tested with 8 clients, despite this not being the optimal configuration, to better simulate real-world FL scenarios. For the UNSW-NB15 dataset, the number of clients is fixed at 12, as this configuration minimizes communication rounds. The fraction fit is then varied to 0.1, 0.5, and 1 to assess its impact.
Increasing the fraction fit effectively reduces communication cycles while maintaining target performance levels. This trend is observed in the Ton-IoT and NSL-KDD datasets, as illustrated in Tables 4(a) and 4(c). However, for the UNSW-NB15 and WUSTL-EHMS datasets, the lowest number of communication rounds is achieved with a fraction fit of 0.5, with a slight increase observed at a fraction fit of 1, as shown in Tables 4(b) and 4(d). This confirms the general trend of reduced communication rounds with higher fraction fit values.
Finally, the analysis investigates the impact of the number of local epochs on performance and communication cycles. The fraction fit is set to 1, as this configuration has been shown to minimize communication rounds. The number of local epochs is varied to 1, 2, 5, and 8.
@@
A notable finding is the consistent reduction in communication cycles with an increase in local epochs while maintaining strong performance. This trend is observed across most datasets, as shown in Tables 5(a), 5(b), and 5(d). However, the NSL-KDD dataset achieves the optimal number of communication rounds with only 2 epochs, as illustrated in Table 5(c), though the results remain close to those obtained with 8 epochs. This further supports the overall trend of reduced communication rounds with increased local epochs.
@@ 4.6.2 XAI results
Previous test results demonstrate the IDS system’s commendable performance in anomaly detection. However, an investigation into why the proposed solution predicts as it does is imperative. To address this concern, the SHAP method is employed, enabling the identification of feature relevance in anomaly detection. The outcomes of the SHAP method can be depicted graphically, utilizing variable-length bars and color coding called beeswarm plots. This visualization effectively demonstrates how each level or range of values of a particular feature positively or negatively influences the classification result.
Each point on the graph represents a feature value. Red points denote higher feature values, while blue points indicate lower feature values. Values on the left side of the x-axis tend toward the normal class, whereas those on the right side tend toward the anomaly class [63]. There are also other types of graphs, such as bar plots. In these plots, the x-axis represents the Shapley value, and the y-axis represents the feature names. Features with the most significant impact are positioned at the top of the graph, while those with the least impact are at the bottom [64].
@@
In this approach, the SHAP value is applied to the final FL model, achieving the best performance using the test dataset. This process produces two types of graphics: beeswarm plots and bar plots.
@@
The SHAP results for the UNSW-NB15 dataset are depicted in Figure 4, while the key features exerting the most significant impact on the anomaly detection classification process, as shown in Figure 4(b), include sttl (Time to Live from the source to the destination), ct_state_ttl (Connection state value of Time to Live), and dttl (Time to Live from the destination to the source). Notably, high values of these features, all linked to Time To Live (TTL), play a pivotal role in anomaly detection, visually represented in the accompanying Figure 4(a).
Prominent patterns emerge, underscoring that elevated values in sttl, ct_state_ttl, and dttl serve as indicators of anomalies, especially concerning TTL. Under normal circumstances, TTL values tend to exhibit stability or fall within specific ranges. However, higher values may signify abnormal extensions of connections, potentially employed to circumvent temporary security mechanisms. It is essential to note that certain attacks comprise very few samples, posing a challenge for ML in discerning crucial features for the detection of such anomalies.
@@
In the ToN-IoT dataset, as illustrated in Figure 5, the features wielding the most significant impact on anomaly detection classification include proto _udp (indicating the use of UDP), dst_port (representing destination ports), conn_state_oth (denoting other unspecified connection states), service_dns (reflecting the use of DNS services), conn_state_SF (indicating an established connection with successful data exchange), conn_state_rej (signifying connection rejection or inability to establish), and src_port (depicting source ports from the endpoint’s TCP/UDP ports), as shown in Figure 5(b). The absence of UDP protocol usage, as demonstrated in Figure 5(a), is pivotal in detecting anomalies, emphasizing its role in discerning abnormal network behavior. Notably, lower destination port numbers indicate ongoing attacks, with attackers often targeting well-known services associated with such ports (e.g., FTP=21, SSH=22, or HTTP=80). Correlated attributes, namely conn_state_oth, conn_state_SF, and conn_state_rej, play a critical role in attack detection. The absence of a specific connection state in conn_state_oth, as depicted in Figure 5(a), suggests potential interception or falsification of traffic, while conn_state_SF points to established connections without successful data exchange, possibly indicating suspicious activities. Simultaneously, conn_state_rej highlights repeated or massive attempts to establish rejected connections, hinting at DDoS attacks or vulnerability exploitation. The absence of DNS port usage, as shown in Figure 5(a), indicates a potential threat, as attackers may manipulate DNS requests to mask their activities. Additionally, elevated values of the source port attribute enhance attack detection, as attackers often initiate attacks from dynamic ports not assigned to well-known services like FTP.
@@
The SHAP results for the WUSTL-EHMS dataset, as illustrated in Figure 7, highlight features influencing the classification process in anomaly detection. These features, outlined in Figure 7(b), include Dintpkt (Destination Inter Packet), dstjitter (Destination Jitter), dstload (Destination Load), load, and srcload (Source Load). Each of these features holds distinct relevance for understanding and characterizing network behavior, especially in the context of health-related host data. Significantly, features associated with health data have a profound impact on detecting attacks, particularly those related to data injection. Anomalies in health-related metrics can be identified by values deviating from typical ranges. For example, elevated pulse_rate or unusually low temperatures, as depicted in Figures 7(a), can signal abnormal conditions. Elevated values of Dintpkt or dstjitter, which are correlated features, may indicate unusual intervals or abnormal temporal variations between packets or data received by the destination. In the context of spoofing, such variations could signal an attempt to manipulate network traffic, concealing or altering the true origin of packets. This observation aligns with the understanding that packets in spoofing scenarios are often manually modified by attackers, sent individually with extended time intervals between them, rather than in an organized flow.
@@
Low values of dstload, load, and srcload also play a significant role in anomaly detection. A destination with low load may be unusual in a normal context, potentially indicating an attacker not targeting system availability but engaging in other types of attacks, such as injection or scanning attacks. This behavior might be explained by the attacker’s intent to remain inconspicuous and avoid detection, highlighting the multifaceted role these features play in discerning anomalies within network behavior.